# Emmanuel Letellier Université Paris Cité, IMJ-PRG CNRS UMR 7586

### v2

@@ 4.2 Characteristic functions of orbital complexes
In what follows, $X$ is an algebraic variety over $\overline{\mathbb{F}}_{q}$ equipped with a geometric Frobenius $F:X\rightarrow X$ (or equivalently, an $\mathbb{F}_{q}$-scheme $X_{o}$ such that $X=X_{o}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$), for more details see for instance [9, Chapter 4].
@@
If $\mathcal{F}$ is the pullback of a complex $\mathcal{F}_{o}$ on $X_{o}$, then it admits a canonical $F$-equivariant structure, see for instance [14, Chapter 1].
@@
The function $\operatorname{{\bf X}}_{\mathcal{F},\varphi}$ does depend on the choice of the isomorphism $\varphi$. However, in all the cases of relevance for this article, we can make a canonical choice of the isomorphism $\varphi$ and we will drop it from the notation. In particular, if $X$ is an algebraic group, we will always assume that $\varphi_{e}=Id$.
@@
[17, Proposition 4.4.13] The set $\{{\bf X}_{IC^{\bullet}_{\overline{C},\mathcal{E}},\varphi}\}$, where $(C,\mathcal{E})$ runs over the pairs as above, forms a basis of the space $\mathcal{C}(G^{F})$ of class functions $G^{F}/G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$.
@@ 4.3 Character-sheaves on tori
[25, Proposition 2.3.1] The map $\mathcal{E}\to\operatorname{{\bf X}}_{\mathcal{E}}$ defines a bijection between the set of $F$-stable isomorphism classes of Kummer local systems on $T$ and the set $\widehat{T^{F}}$ of linear characters of $T^{F}$.
@@
Consider $T=\mathbb{G}_{m}$ with the canonical Frobenius $F(x)=x^{q}$ for $x\in\mathbb{G}_{m}$. In this case, we have $T^{F}=\mathbb{F}_{q}^{*}$. Consider a linear character $\alpha:\mathbb{F}_{q}^{*}\to\mathbb{C}^{*}$ and let $n$ be the order of $\alpha$. In particular, $n$ divides $q-1$. Fix a surjection $q_{n}:\mathbb{F}_{q}^{*}\to\mathbb{Z}/n\mathbb{Z}$ (by sending a generator $\zeta$ of the cyclic group $\mathbb{F}_{q}^{*}$ to its subgroup of order $n$ generated by $\zeta^{\frac{q-1}{n}}$). Since $\alpha^{n}=1$, there exists a linear character $\mu:\mathbb{Z}/n\mathbb{Z}\to\mathbb{C}^{*}$ such that $\mu\circ q_{n}=\alpha$.
@@ 4.4 Grothendieck-Springer resolution and geometric induction
For more details and references on this section, see [15, §6.1, 6.2] and also [16, §2.4, 2.7].
@@
Consider $\mathcal{F}\in D_{c}^{b}(T)$ equipped with both a $W$-equivariant structure $\theta$ and an $F$-equivariant structure $\varphi:F^{*}(\mathcal{F})\simeq\mathcal{F}$ such that the two are compatible, namely the following diagram commutes for all $w\in W$, see [15, Diagram (6.1)]:
@@
For any $\chi\in\widehat{W}$, we see from this compatibility condition (see [15, Remark 6.1.3] for more details) that the canonical $F$-equivariant structure $\varphi^{G}:F^{*}\left(\mathcal{I}_{T}^{G}(\mathcal{F})\right)\simeq\mathcal{I}_{T}^{G}(\mathcal{F})$ restricts to an $F$-equivariant structure $\varphi^{G}_{\chi}:F^{*}(\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi})\simeq\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ and
@@ 4.5 Character-sheaves on $\mathrm{SL}_{2}$
Thanks to the above theorem, the characteristic functions of the $F$-stable character-sheaves are closely related to the irreducible characters of $G^{F}$. When $G=\operatorname{GL}_{n}$, the characteristic functions of the $F$-stable character-sheaves are exactly (up to a sign) the irreducible characters of $\operatorname{GL}_{n}(\mathbb{F}_{q})$. This is not true in general (as we will see in details with $\operatorname{SL}_{2}$). Notice that there is in fact no reason that the decomposition of $\mathcal{I}_{T}^{G}(\mathcal{L})$ as a direct sum of simple objects in the category of perverse sheaves would agree with the decomposition of the character $R_{T}^{G}({\bf X}_{\mathcal{L}})$ as a direct sum of irreducible characters.
@@
Fix a an $F$-stable Kummer local system $\mathcal{L}$.
@@
Notice that if $\mathcal{L}$ is $W$-equivariant on $T=T_{1}$, then $\mathcal{L}$ will be also $\sigma F$-stable. Since the pairs $(T_{1},\sigma F)$ is $G$-conjugate to $(T_{\sigma},F)$, we get an $F$-stable Kummer local system $\mathcal{L}_{\sigma}$ on $T_{\sigma}$ and we have
@@
$(f_{2})_{*}(\overline{{\mathbb{Q}}}_{\ell})=\overline{{\mathbb{Q}}}_{\ell}\oplus\mathcal{A}_{o}.$
It corresponds to the character $\alpha_{o}$ of $\mathbb{F}_{q}^{*}\simeq T_{1}^{F}$ (which take the value $1$ at squares and $-1$ at non-squares). Since $f_{2}$ is $W$-equivariant, the local system $\mathcal{A}_{o}$ must be $W$-equivariant and by (4.4.2) we have
@@ 5.2 Some result for $\mathrm{GL}_{2}$
We have the following definition [19, Definition 6.8.6].
@@
[19, Theorem 6.10.1] If $R$ is generic then
@@ 5.3 From $\mathrm{GL}_{2}$ to $\mathrm{SL}_{2}$
$R_{i}=\begin{cases}R_{T^{\prime}_{1}}^{\operatorname{GL}_{2}}(1,\gamma_{i})&\text{ if }1\leq i\leq k_{1},\\ -R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega_{i})&\text{ if }k_{1}+1\leq i\leq k_{2}.\end{cases}$
@@ 5.4 Generic $k$-tuples of semisimple characters of $\mathrm{SL}_{2}(\mathbb{F}_{q})$
$R_{T}^{\operatorname{GL}_{2}}(\theta)\otimes(\lambda\circ\det)=R_{T}^{Gl_{2}}(\theta\otimes(\lambda\circ\det)).$
@@
Notice that eq. (5.4.1) above holds if and only if
@@
i.e. eq(5.4.1) is equivalent to the Condition $(2)$ of Definition 5.4.1 for the $k$-tuple $R^{{}^{\prime}}$ and the elements $w_{1},\dots,w_{k}$.

### v3

@@ 4.2 Characteristic functions of orbital complexes
In what follows, $X$ is an algebraic variety over $\overline{\mathbb{F}}_{q}$ equipped with a geometric Frobenius $F:X\rightarrow X$ (or equivalently, an $\mathbb{F}_{q}$-scheme $X_{o}$ such that $X=X_{o}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$), for more details see for instance [8, Chapter 4].
@@
If $\mathcal{F}$ is the pullback of a complex $\mathcal{F}_{o}$ on $X_{o}$, then it admits a canonical $F$-equivariant structure, see for instance [13, Chapter 1].
@@
The function $\operatorname{{\bf X}}_{\mathcal{F},\varphi}$ does depend on the choice of the isomorphism $\varphi$. However, in all the cases of relevance for this article, we can make a canonical choice of the isomorphism $\varphi$ and we will often drop it from the notation. In particular, if $X$ is an algebraic group, we will always assume that $\varphi_{e}=Id$.
@@
[16, Proposition 4.4.13] The set $\{{\bf X}_{IC^{\bullet}_{\overline{C},\mathcal{E}},\varphi}\}$, where $(C,\mathcal{E})$ runs over the pairs as above, forms a basis of the space $\mathcal{C}(G^{F})$ of class functions $G^{F}/G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$.
@@ 4.3 Character-sheaves on tori
[24, Proposition 2.3.1] The map $\mathcal{E}\to\operatorname{{\bf X}}_{\mathcal{E}}$ is an isomorphism between the group of $F$-stable isomorphism classes of Kummer local systems on $T$ and the group $\widehat{T^{F}}$ of linear characters of $T^{F}$.
@@
Consider $T=\mathbb{G}_{m}$ with the Frobenius $F(x)=x^{q}$ for $x\in\mathbb{G}_{m}$. In this case, we have $T^{F}=\mathbb{F}_{q}^{*}$. Consider a linear character $\alpha:\mathbb{F}_{q}^{*}\to\mathbb{C}^{*}$ and let $n$ be the order of $\alpha$. In particular, $n$ divides $q-1$. Fix a surjection $q_{n}:\mathbb{F}_{q}^{*}\to\mathbb{Z}/n\mathbb{Z}$ (by sending a generator $\zeta$ of the cyclic group $\mathbb{F}_{q}^{*}$ to its subgroup of order $n$ generated by $\zeta^{\frac{q-1}{n}}$). Since $\alpha^{n}=1$, there exists a linear character $\mu:\mathbb{Z}/n\mathbb{Z}\to\mathbb{C}^{*}$ such that $\mu\circ q_{n}=\alpha$.
@@ 4.4 Grothendieck-Springer resolution and geometric induction
For more details and references on this section, see [14, §6.1, 6.2] and also [15, §2.4, 2.7].
@@
Consider $\mathcal{F}\in D_{c}^{b}(T)$ equipped with both a $W$-equivariant structure $\theta$ and an $F$-equivariant structure $\varphi:F^{*}(\mathcal{F})\simeq\mathcal{F}$ such that the two are compatible, namely the following diagram commutes for all $w\in W$, see [14, Diagram (6.1)]:
@@
For any $\chi\in\widehat{W}$, we see from this compatibility condition (see [14, Remark 6.1.3] for more details) that the canonical $F$-equivariant structure $\varphi^{G}:F^{*}\left(\mathcal{I}_{T}^{G}(\mathcal{F})\right)\simeq\mathcal{I}_{T}^{G}(\mathcal{F})$ restricts to an $F$-equivariant structure $\varphi^{G}_{\chi}:F^{*}(\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi})\simeq\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ and
@@ 4.5 Character-sheaves on $\mathrm{SL}_{2}$
Thanks to the above theorem, the characteristic functions of the $F$-stable character-sheaves are closely related to the irreducible characters of $G^{F}$. When $G=\operatorname{GL}_{n}$, the characteristic functions of the $F$-stable character-sheaves are exactly (up to a sign) the irreducible characters of $\operatorname{GL}_{n}(\mathbb{F}_{q})$. This is not true in general (as we will see in details with $\operatorname{SL}_{2}$).
Notice that there is in fact no reason that the decomposition of $\mathcal{I}_{T}^{G}(\mathcal{L})$ as a direct sum of simple objects in the category of perverse sheaves would agree with the decomposition of the character $R_{T}^{G}({\bf X}_{\mathcal{L}})$ as a direct sum of irreducible characters.
@@
Fix a an $F$-stable Kummer local system $\mathcal{L}$ on $T$.
@@
Notice that if $\mathcal{L}$ is $W$-equivariant on $T=T_{1}$, then $\mathcal{L}$ will be also $\sigma F$-stable. Since the pairs $(T_{1},\sigma F)$ is $G$-conjugate to $(T_{\sigma},F)$ (see Diagram (4.1.1)), we get an $F$-stable Kummer local system $\mathcal{L}_{\sigma}$ on $T_{\sigma}$ and we have
@@
$(f_{2})_{*}(\overline{{\mathbb{Q}}}_{\ell})=\overline{{\mathbb{Q}}}_{\ell}\oplus\mathcal{A}_{o}$
where $f_{2}:T_{1}\rightarrow T_{1}$, $z\mapsto z^{2}$.
The Kummer local system $\mathcal{A}_{o}$ corresponds to the character $\alpha_{o}$ of $\mathbb{F}_{q}^{*}\simeq T_{1}^{F}$ (which take the value $1$ at squares and $-1$ at non-squares). Since $f_{2}$ is $W$-equivariant, the local system $\mathcal{A}_{o}$ must be $W$-equivariant and by (4.4.2) we have
@@ 5.2 Some result for $\mathrm{GL}_{2}$
We have the following definition [18, Definition 6.8.6].
@@
[18, Theorem 6.10.1] If $R$ is generic then
@@ 5.3 From $\mathrm{GL}_{2}$ to $\mathrm{SL}_{2}$
$R_{i}=\begin{cases}R_{T_{1}}^{\operatorname{GL}_{2}}(1,\gamma_{i})&\text{ if }1\leq i\leq k_{1},\\ -R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega_{i})&\text{ if }k_{1}+1\leq i\leq k_{2}.\end{cases}$
@@ 5.4 Generic $k$-tuples of semisimple characters of $\mathrm{SL}_{2}(\mathbb{F}_{q})$
$R_{T}^{\operatorname{GL}_{2}}(\theta)\otimes(\lambda\circ\det)=R_{T}^{\operatorname{GL}_{2}}(\theta\otimes(\lambda\circ\det)).$
@@
Notice that (5.4.1) above holds if and only if
@@
i.e. (5.4.1) is equivalent to the Condition $(2)$ of Definition 5.4.1 for the $k$-tuple $R^{{}^{\prime}}$ and the elements $w_{1},\dots,w_{k}$.